OpenStudy (anonymous):

please help. What is the value of x? x/4-5=6

OpenStudy (anonymous):

@Stormtrooper can you help please?

OpenStudy (anonymous):

Okay x/4-5=6 Add 5 to both sides. x/4=11 Multiply four to both sides x=44 Because you canceled out everything else you now have your single variable

OpenStudy (denonakavro):

x/4-5 = 6 add 5 both sides x/4 = 11 multiply 4 both sides x = 44

OpenStudy (anonymous):

ok thx again i might ask another question later

OpenStudy (anonymous):

No problem.